OpenStudy (anonymous):

Find the slope of line passing through the points (5, 9) and (5, -2)

OpenStudy (anonymous):

y2-y1/x2-x1=slope formula so the slope is 0/-11 so the slope is just 0

OpenStudy (anonymous):

Which means it is a horizontal line. If the slope is over 0, it would be undefined and be a vertical line.

OpenStudy (anonymous):

there is no slope .....cuz y2-y1/x2-x1=-2-9/5-5=-11/0...so it's undefined so there is no slope

OpenStudy (anonymous):

oh, HERR DERR yeah, i wrote it right but solved it wrong ;)